OpenStudy (anonymous):

Can someone help me simplify this. tan^2(x)-sin^2(x)

OpenStudy (anonymous):

\[ \tan^2x-\sin^2x = \frac{\sin^2x}{\cos^2x}-\frac{\sin^2x}{1} \]

OpenStudy (anonymous):

so what do you do next

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Do you know how to subtract fractions?

OpenStudy (anonymous):